# FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models

### v1

@@ Abstract
Recent research in federated large language models (LLMs) has primarily focused on enabling clients to fine-tune their locally deployed homogeneous LLMs collaboratively or on transferring knowledge from server-based LLMs to small language models (SLMs) at downstream clients. However, a significant gap remains in the simultaneous mutual enhancement of both the server’s LLM and clients’ SLMs. To bridge this gap, we propose FedMKT, a parameter-efficient federated mutual knowledge transfer framework for large and small language models. This framework is designed to adaptively transfer knowledge from the server’s LLM to clients’ SLMs while concurrently enriching the LLM with clients’ unique domain insights. We facilitate token alignment using minimum edit distance (MinED) and then selective mutual knowledge transfer between client-side SLMs and a server-side LLM, aiming to collectively enhance their performance. Through extensive experiments across three distinct scenarios, heterogeneous, homogeneous, and one-to-one, we evaluate the effectiveness of FedMKT using various public LLMs and SLMs on a range of NLP text generation tasks. Empirical results demonstrate significant performance improvements in clients’ SLMs with the aid of the LLM. Furthermore, the LLM optimized by FedMKT achieves a performance comparable to that achieved through direct fine-tuning based on clients’ data, highlighting the effectiveness and adaptability of FedMKT.
@@ 1 Introduction
Large Language Models (LLMs) have emerged as a transformative force in artificial intelligence, profoundly altering our perception of natural language processing capabilities. The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022) and GPT-4 OpenAI (2023), with their billions of parameters, has sparked the imagination of both researchers and practitioners, owing to their exceptional performance across diverse text generation tasks. This trend signifies a growing emphasis on developing and utilizing large-scale pre-train models, underscoring their increasing importance in advancing the frontiers of AI. The remarkable success of LLMs in general NLP tasks has also fueled the release of numerous high-performing open-source LLMs, including LLaMA Touvron et al. (2023), ChatGLM Du et al. (2022), OPT Zhang et al. (2022a), igniting a wave of enthusiasm for further exploration and application.
Despite their widespread success in various general NLP tasks, LLMs face challenges that hinder their adoption in domain-specific applications Kang et al. (2023) Fan et al. (2023). The primary challenges include:
  • Domain-Specific Knowledge Privacy. Downstream clients are unable to access the LLM parameters when the LLM is black-box. Hence, they must transmit their domain-specific data to the server that owns the LLM owners for fine-tuning purposes. This process inevitably compromises the privacy of clients’ sensitive, domain-specific data.
  • Constrained Computing Resources. Downstream clients may face limitations in computing resources, which can hinder their ability to fine-tune downloaded LLMs, even if they have access to the server’s LLM. The fine-tuning process requires substantial computational and storage resources, posing a significant barrier for small and medium-sized companies with limited resources to adopt LLMs. Consequently, these companies, which act as clients in FL, are often compelled to fine-tune smaller language models (SLMs) using their domain-specific data, which may result in inferior model performance.
  • Mutual Knowledge Transfer Between the LLM and SLMs. The optimization of both the LLM and SLMs across different parties establishes a positive feedback loop, enabling the continuous evolution of both the LLM and SLMs. Initially, the LLM on the server can disseminate general knowledge and capabilities to the clients’ SLMs. Subsequently, these domain-specific SLMs can return their industry-specific knowledge to the server’s LLM, fostering a collaborative learning environment. However, a significant challenge arises from the inherent model heterogeneity between the LLM and SLMs, particularly when aligning distributions of output logits. The mismatch between the tokenizers of different LLM and SLMs poses a notable obstacle. Furthermore, the mutual transfer of knowledge between the server’s LLM and clients’ SLMs remains a largely unexplored area in academic literature, warranting further investigation.
To address these challenges, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to enhance the performance of both large and small language models. By leveraging the complementary strengths of federated learning and knowledge distillation techniques, FedMKT facilitates effective mutual knowledge transfer between clients’ small language models (SLMs) and a large language model (LLM) located at the server.
@@
Extensive experiments conducted on various public LLM and SLMs, including GPT-2 Radford et al. (2019), OPT Zhang et al. (2022a), Bloom Scao et al. (2022), LLaMa Touvron et al. (2023), demonstrate the competitive performance of our FedMKT framework across a range of NLP text generation tasks. In our comprehensive evaluation, we test our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One. The results show that the SLMs can achieve significant enhancements with the support of the LLM, while the LLM can deliver comparable results to fine-tuning with all clients’ domain data directly. Overall, FedMKT represents a practical and effective solution for grounding LLMs in real-world applications while maintaining data privacy protections.
@@
  • Federated Mutual Knowledge Transfer Framework. FedMKT introduces a novel federated mutual knowledge transfer framework that enables effective knowledge transfer between large language models (LLMs) deployed on the server and small language models (SLMs) residing on clients. This framework fills the gap by simultaneously enhancing both the server’s LLM and the clients’ SLMs.
@@
  • Empirical Evaluation and Performance Enhancement. Through extensive experiments on various public LLMs and SLMs across NLP text generation tasks, FedMKT demonstrates significant performance enhancements in clients’ SLMs with the aid of the server’s LLM. Additionally, the LLM optimized by FedMKT achieves comparable performance to direct fine-tuning on clients’ data, highlighting the effectiveness and adaptability of the proposed framework for grounding LLMs in real-world applications while maintaining data privacy protections.
Figure 1: FedMKT (Federated Mutual Knowledge Transfer for Large and Small Language Models Framework.)
@@ 2.1 Model Heterogeneous Federated Learning
Federated learning has emerged as a promising approach for collaborative machine learning across multiple devices while preserving data privacy. However, traditional federated learning frameworks often assume homogeneity among client models, which may not hold in real-world scenarios due to varying computational capabilities, network conditions, and data distributions. This limitation has led to the development of model heterogeneous federated learning (MHFL), which aims to address the challenges associated with heterogeneity. Initial research in MHFL primarily concentrated on addressing heterogeneity in model architectures. Various approaches have been proposed to accommodate clients with different model architectures participating in a federated learning task. These methods typically involve techniques such as knowledge distillation Hinton et al. (2015), mutual learning and split learning that can handle heterogeneous models.
Knowledge distillation-based MHFL methods, such as FedMD Li and Wang (2019) and FedET Cho et al. (2022), involve the server aggregating the output logits of different clients’ heterogeneous models on a public dataset to construct global logits. Each client then calculates the distance between the global logits and the local logits of their local heterogeneous model on the same public data sample, using this as a distillation loss for training the local model.
Mutual learning-based MHFL, such as Deep Mutual Learning (DML) Zhang et al. (2018), PFML Yang et al. (2021) and FedLoRA Yi et al. (2023), design a small homogeneous model and a large heterogeneous model in each client. These methods integrate mutual learning into personalized federated learning, aiming to enhance the performance of both the global and personalized heterogeneous models.
Split learning-based MHFL approaches, such as FedClassAvg Jang et al. (2022) and CHFL Liu et al. (2022), share a homogeneous classifier to improve model classification while personalizing the local feature extractor.
While previous work has mainly focused on computer vision (CV) scenarios, there has been limited exploration of MHFL in large language model (LLM). This gap in the literature motivates our work, which aims to explore MHFL in the context of LLMs.
@@ 2.2 Federated Learning for LLMs
Adapting large language models (LLMs) Yang et al. (2023); Zhou et al. (2023); OpenAI (2022, 2023); Touvron et al. (2023) to downstream tasks typically involve the comprehensive fine-tuning of all model parameters, a process known as full fine-tuning. However, this approach can be prohibitively expensive, particularly for domain-specific tasks. To address this challenge, Parameter-Efficient Fine-Tuning (PEFT) methods Houlsby et al. (2019); He et al. (2021); Lester et al. (2021); Li and Liang (2021); Hu et al. (2021) have been proposed. These methods efficiently adapt LLMs to specific domains or tasks by fine-tuning only a subset of model parameters while keeping the rest frozen.
PEFT methods offer a direct solution to the issues of communication overhead and fine-tuning costs in federated learning (FL) for large language models (LLMs). A number of studies have built upon PEFT methods in the context of FL for LLMs, including FedPETuning Zhang et al. (2022b), Federated Adapter Tuning Cai et al. (2022), Federated Prompt Tuning Zhao et al. (2022), and FATE-LLM Fan et al. (2023). Specifically, the FedPETuning methods proposed by Zhang et al. (2022b) have demonstrated a significant reduction in communication overhead, achieving a reduction of 1 to 2 orders of magnitude compared to full fine-tuning in the FL setting. Additionally, they found that PETuning methods can effectively reduce local model adaptation costs for clients in FL systems. These findings imply that FL clients, such as devices with limited storage capacity, can greatly benefit from PEFT methods. These methods enable the sharing of LLMs across different tasks while maintaining only a few parameters for each task, thereby reducing the storage requirement. By leveraging PEFT methods, FL clients can efficiently adapt LLMs to their specific needs while minimizing communication overhead and fine-tuning costs.
3 The Proposed Method
In this section, we introduce FedMKT, an innovative and parameter-efficient federated mutual knowledge transfer for large and small language models framework. This framework encompasses fine-tuning scenarios in an FL setting where an LLM is run on the server while SLMs are distributed among the clients. A notable aspect of this architecture is the heterogeneity in designs among the client SLMs and their contrast with the server LLM. We begin by defining the specific problem addressed in this study, followed by a detailed introduction to our approach. Finally, we delve into the computational and communication complexities and the privacy-preserving aspects of our FedMKT framework.
@@ 3.1 Problem Definition
We consider the federated learning setting, involving one server that owns an LLM $f_{\psi}$ parameterized by $\psi$ and $K$ clients that each client $k$ has an SLM $g_{\phi_{k}}$ parameterized by ${\phi}_{k}$. We assume that the $K$ clients execute the same text generation task, but they hold heterogeneous SLM $g_{\phi_{k}}$.
The server and clients aim to collaboratively enhance the performance of the LLM and SLMs through federated learning without ever disclosing any private data. The threat model we consider is a standard semi-honest security model, wherein a semi-honest server(S) attempts to deduce the private data belonging to the client(c). Specifically,
  • Each client possesses its own local private dataset $\mathcal{D}^{k}$ and owns a common public dataset $D^{p}$. Each client $k$ aims to collectively train its local private SLM $g_{\phi_{k}}$ without divulging their private data. Subsequently, each client computes the output logits on the public dataset and securely uploads them to the server. The objective can be formulated as follows:
    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{1}(\phi_{1},\phi_{2},...,\phi_{K};\{\mathcal{D}^{k}\}_{k=1}^{K})$ (1)
  • The server also owns the common public dataset $D^{p}$, and receives the output logits of different clients’ SLMs on the public dataset. Subsequently, the server computes the distillation loss by comparing these client logits with the output logits produced by its own LLM on the public dataset. The objective can be formulated as follows:
    $\min_{\psi}\mathcal{L}_{2}(\psi;\mathcal{D}^{p},{\phi_{1},\phi_{2},...,\phi_{K}})$ (2)
@@
  • Once the knowledge transfer from SLMs to LLM is completed on the server, the server dispatches the LLM’s output logits on the public dataset $D^{p}$ to all the clients. Subsequently, the clients compute the distillation loss by comparing LLM output logits with SLMs’ output logits on the public dataset $D^{p}$. The objective can be formulated as follows:
    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{3}(\phi_{1},\phi_{2},...,\phi_{K};\mathcal{D}^{p},{\psi})$ (3)
    The clients aim to transfer knowledge from LLM $f_{\psi}$ to SLMs $g_{\phi_{k}}$ to enhance SLMs in the clients.
FedMKT solves the optimization problems formulated in Eq.(1), Eq.(2), and Eq.(3) in an efficient and privacy-preserving manner. We elaborate on FedMKT in Section 3.2.
3.2 FedMKT
FedMKT represents an innovative and parameter-efficient framework for federated mutual knowledge transfer between large and small language models. Its primary objective is to foster the collaborative advancement of both server-side LLM and client-side SLMs. Additionally, in our work, we specifically focus on text-generation tasks. We illustrate the FedMKT in Figure 1 and describe the associated training algorithm in Algorithm 1.
@@ 3.2 FedMKT
3:$T$: total number of rounds;
4:$\theta^{0}_{k}$: the initialized global adapter of SLM $g_{\phi_{k}+\theta_{k}}$;
5:$\omega^{0}$: the initialized adapter of LLM $f_{\psi+\omega}$;
6:$\eta_{\omega}$: the learning rate for optimizing LLM $f_{\psi+\omega}$;
7:$\eta_{\theta}$: the learning rate for optimizing SLM $g_{\phi_{k}+\theta_{k}}$.
@@
9:for $t$ in communication round $T$ do
@@
11:     $p_{k}^{t},l_{k}^{t}\leftarrow\textbf{ClientUpdate1}(t)$
12:     $\triangleright$ selective knowledge transfer from SLMs to LLM based on $\mathcal{D}^{p}$
13:     Token Alignment from SLMs to LLM on $\mathcal{D}^{p}$.
14:     for each epoch $r\in[R]$ do
15:         $\omega^{t,r+1}\leftarrow\omega^{t,r}-\eta_{\omega}\bigtriangledown\mathcal{L}_{2}$
16:     end for
17:     $\omega^{t+1}=\omega^{t,R}$
18:     $p^{t}\leftarrow ComputeLogits(\mathcal{D}^{p})$
19:     $l^{t}\leftarrow ComputeCELoss(\mathcal{D}^{p})$
@@
21:
@@
23:     for each client $k$ (in parallel) do
24:         $\triangleright$ local fine-tuning based on $\mathcal{D}^{k}$
25:         for each local epoch $e\in[E]$ do
26:              $\theta_{k}^{t,e+1}\leftarrow\theta_{k}^{t,e}-\eta_{\theta}\bigtriangledown\ell_{\text{TA}}^{k}$
27:         end for
28:         $p_{k}^{t}\leftarrow ComputeLogits(\mathcal{D}^{p})$
29:         $l_{k}^{t}\leftarrow ComputeCELoss(\mathcal{D}^{p})$
30:     end for
31:
32:     //ClientUpdate2($p^{t},l^{t}$):
33:     for each client $k$ (in parallel) do
34:         Token Alignment from LLM to SLM $g_{\phi_{k}+\theta_{k}}$ on $\mathcal{D}^{p}$.
35:         $\triangleright$ selective knowledge transfer from LLM to SLMs based on $\mathcal{D}^{p}$
36:         for each local epoch $e\in[E,2E]$ do
37:              $\theta^{t,e+1}_{k}\leftarrow\theta^{t,e}_{k}-\eta_{\theta}\bigtriangledown\mathcal{L}_{3}$
38:         end for
39:         $\theta^{t+1}_{k}=\theta^{t,2E}_{k}$
40:     end for
41:end for

### v2

@@ Abstract
Recent research in federated large language models (LLMs) has primarily focused on enabling clients to fine-tune their locally deployed homogeneous LLMs collaboratively or on transferring knowledge from server-based LLMs to small language models (SLMs) at downstream clients. However, a significant gap remains in the simultaneous mutual enhancement of both the server’s LLM and clients’ SLMs. To bridge this gap, we propose FedMKT, a parameter-efficient federated mutual knowledge transfer framework for large and small language models. This framework is designed to adaptively transfer knowledge from the server’s LLM to clients’ SLMs while concurrently enriching the LLM with clients’ unique domain insights. We facilitate token alignment using minimum edit distance (MinED) and then selective mutual knowledge transfer between client-side SLMs and a server-side LLM, aiming to collectively enhance their performance. Through extensive experiments across three distinct scenarios, we evaluate the effectiveness of FedMKT using various public LLMs and SLMs on a range of NLP text generation tasks. Empirical results demonstrate that FedMKT simultaneously boosts the performance of both LLMs and SLMs.
FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models
Tao Fan1, 2, Guoqiang Ma2, Yan Kang2, Hanlin Gu2, Yuanfeng Song2, Lixin Fan2, Kai Chen1, Qiang Yang1,2 1 Hong Kong University of Science and Technology, China 2 WeBank, China Correspondence: tfanac@cse.ust.hk
@@ 1 Introduction
Large Language Models (LLMs) have emerged as a transformative force in artificial intelligence, profoundly altering our perception of natural language processing capabilities. The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022), and LLaMA Touvron et al. (2023) with their billions of parameters, has sparked the imagination of both researchers and practitioners, owing to their exceptional performance across diverse text generation tasks. Despite their widespread success in various general NLP tasks, LLMs face challenges that hinder their adoption in domain-specific applications Kang et al. (2023) Fan et al. (2023). The primary challenges include domain-specific knowledge Privacy, constrained computing resources, and mutual knowledge transfer between the LLM and SLMs. A significant challenge arises from the inherent model heterogeneity between the LLM and SLMs, particularly when aligning distributions of output logits. The mismatch between the tokenizers of different LLM and SLMs poses a notable obstacle. Furthermore, the mutual transfer of knowledge between the server’s LLM and clients’ SLMs remains a largely unexplored area in academic literature, warranting further investigation.
To fill these gaps, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to enhance the performance of both large and small language models. By leveraging the complementary strengths of federated learning and knowledge distillation, FedMKT facilitates effective mutual knowledge transfer between clients’ SLMs and the LLM owned by the server.
@@
  • Federated Mutual Knowledge Transfer Framework. FedMKT introduces a novel federated mutual knowledge transfer framework that enables effective knowledge transfer between an LLM deployed on the server and SLMs residing on clients. This framework fills the gap by simultaneously enhancing both the server’s LLM and the clients’ SLMs.
@@
  • Empirical Evaluation and Performance Enhancement. Extensive experiments conducted based on various publicly available LLMs and SLMs demonstrate the competitive performance of FedMKT across a wide range of NLP text-generation tasks. We evaluate FedMKT with heterogeneous, Homogeneous, and One-to-One settings. The results show that the performance of SLMs can be significantly enhanced with the help of the LLM, while the LLM can deliver comparable results to fine-tuning with all clients’ data centralized.
Figure 1: Overview of the proposed FedMKT workflow. Each communication round of FedMKT involves 11 steps to fine-tune the server’ LLM and clients’ SLMs.
@@ 2.1 Model Heterogeneous Federated Learning
Model heterogeneous federated learning (MHFL) aims to address the challenges associated with heterogeneity in federated learning. Initial research in MHFL primarily concentrated on addressing heterogeneity in model architectures. Various approaches have been proposed to accommodate clients with different model architectures participating in a federated learning task. These methods typically involve techniques such as knowledge distillation Hinton et al. (2015), mutual learning and split learning that can handle heterogeneous models. Knowledge distillation-based MHFL methods, such as FedMD Li and Wang (2019) and FedET Cho et al. (2022), involve the server aggregating the output logits of different clients’ heterogeneous models on a public dataset to construct global logits. Mutual learning-based MHFL, such as Deep Mutual Learning (DML) Zhang et al. (2018), PFML Yang et al. (2021) and FedLoRA Yi et al. (2023), design a small homogeneous model and a large heterogeneous model in each client. Split learning-based MHFL approaches, such as FedClassAvg Jang et al. (2022) and CHFL Liu et al. (2022), share a homogeneous classifier to improve model classification while personalizing the local feature extractor.
While previous works have mainly focused on computer vision scenarios, the literature has limitedly explored MHFL in LLMs. This gap motivates this study, which aims to explore MHFL in the context of LLMs.
@@ 2.2 Federated Learning for LLMs
Parameter-Efficient Fine-Tuning (PEFT) methods Houlsby et al. (2019); He et al. (2021); Lester et al. (2021); Li and Liang (2021); Hu et al. (2021) offer a direct solution to the issues of communication overhead and fine-tuning costs in federated learning (FL) for LLMs. A number of studies have built upon PEFT methods in the context of FL for LLMs, including FedPETuning Zhang et al. (2022b), Federated Adapter Tuning Cai et al. (2022), Federated Prompt Tuning Zhao et al. (2022), and FATE-LLM Fan et al. (2023). For example, the FedPETuning Zhang et al. (2022b) has demonstrated a significant reduction in communication overhead, reducing 1 to 2 orders of magnitude compared to full fine-tuning in the FL setting. These findings imply that FL clients, such as devices with limited storage capacity, can greatly benefit from PEFT methods. These methods enable the sharing of LLMs across different tasks while maintaining only a few parameters for each task, thereby reducing the storage requirement. By leveraging PEFT methods, FL clients can efficiently adapt LLMs to their specific needs while minimizing communication overhead and fine-tuning costs.
3 The Proposed FedMKT Method
In this section, we introduce FedMKT, an innovative and parameter-efficient federated mutual knowledge transfer approach for large and small language models. The FedMKT primarily comprises two key modules: Bidirectional Token Alignment and Selective Mutual Knowledge Transfer. We will elaborate on these two modules in Section 3.2 and Section 3.3, respectively after we define the problem we try to address in Section 3.1.
@@ 3.1 Problem Definition
We consider the federated learning setting, involving one server that owns an LLM $f_{\psi}$ parameterized by $\psi$ and $K$ clients that each client $k$ has an SLM $g_{\phi_{k}}$ parameterized by ${\phi}_{k}$. Each client owns a local private dataset $\mathcal{D}_{k}$ with $N$ training samples, and all clients and the server share a public dataset $\mathcal{D}_{p}$.
The server and clients aim to collaboratively enhance the performance of the LLM and SLMs through federated learning without disclosing any private data. We assume that the $K$ clients execute the same text generation task, but they may hold heterogeneous or homogeneous SLM models. The collaboration between clients and the server involves the following sub-procedures:
  • Each client $k$ trains its SLM $g_{\phi_{k}}$ using its private data $\mathcal{D}_{k}$. The objective is formulated as follows:
    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{1}(\phi_{1},\phi_{2},...,\phi_{K};\{\mathcal{D}_{k}\}_{k=1}^{K})$ (1)
  • Each client computes the output logits on $\mathcal{D}_{p}$ and securely uploads them to the server. Upon receiving output logits of all clients, the server computes the distillation loss by comparing these client logits with the output logits produced by its own LLM on $\mathcal{D}_{p}$. The objective can be formulated as follows:
    $\min_{\psi}\mathcal{L}_{2}(\psi;\mathcal{D}_{p},{\phi_{1},\phi_{2},...,\phi_{K}})$ (2)
@@
  • The server dispatches the LLM’s output logits on $\mathcal{D}_{p}$ to all the clients. Subsequently, the clients compute the distillation loss by comparing LLM output logits with SLMs’ output logits on $\mathcal{D}_{p}$. The objective can be formulated as follows:
    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{3}(\phi_{1},\phi_{2},...,\phi_{K};\mathcal{D}_{p},{\psi})$ (3)
    The clients aim to transfer knowledge from LLM $f_{\psi}$ to enhance their SLMs.
We consider the server semi-honest, meaning that the server may try to recover the private data of clients from the information it observes.
FedMKT solves the optimization problems formulated in Eq.(1), Eq.(2), and Eq.(3) in an efficient and privacy-preserving manner. We illustrate the workflow of FedMKT in Figure 1 and elaborate on the associated training algorithm in Algorithm 1.
@@
3:$T$: total number of communication rounds;
4:$R$: local number of rounds in the server;
5:$E$: local number of rounds in the client;
6:$\eta_{\omega}$: the learning rate of LLM $f_{\psi+\omega}$;
7:$\eta_{\theta}$: the learning rate of SLM $g_{\phi_{k}+\theta_{k}}$.
@@
9:// Server side:
10:for $t$ in communication round $T$ do
@@
12:     Token Alignment from SLMs to LLM.
13:     $\tilde{\mathcal{S}}_{0}\leftarrow\textbf{DualMinCE}(\mathcal{D}_{p},f_{\psi+\omega},\{\mathcal{S}_{k}\}_{k=1}^{K})$.
14:     // knowledge transfer based on $\mathcal{D}_{p}$ and $\tilde{\mathcal{S}}_{0}$.
15:     for each epoch $r\in[R]$ do
16:         $\omega^{t,r+1}\leftarrow\omega^{t,r}-\eta_{\omega}\bigtriangledown\mathcal{L}_{2}$.
17:     end for
18:     $\omega^{t+1}=\omega^{t,R}$.
19:     Compute $\mathcal{S}_{0}=\{l^{i}_{0},p^{i}_{0}\}_{i=1}^{N}$ based on $\mathcal{D}_{p}$.
@@
21:end for
22:
23:ClientUpdate1($t$):
24:for each client $k$ (in parallel) do
25:     // local fine-tuning based on $\mathcal{D}_{k}$.
26:     for each local epoch $e\in[E]$ do
27:         $\theta_{k}^{t,e+1}\leftarrow\theta_{k}^{t,e}-\eta_{\theta}\bigtriangledown\ell_{\text{TA}}$.
28:     end for
29:     Compute $\mathcal{S}_{k}=\{l^{i}_{k},p^{i}_{k}\}_{i=1}^{N}$ based on $\mathcal{D}_{p}$.
30:end for
31:Upload $\{\mathcal{S}_{k}\}_{k=1}^{K}$ to the server
32:
33:ClientUpdate2($t,\mathcal{S}_{0}$):
34:for each client $k$ (in parallel) do
35:     Token Alignment from LLM to SLMs.
36:     $\tilde{\mathcal{S}}_{k}\leftarrow$ $\textbf{DualMinCE}(\mathcal{D}_{p}$, $g_{\phi_{k}+\theta_{k}}$, $\mathcal{S}_{0}$).
37:     // knowledge transfer based on $\mathcal{D}_{p}$ and $\tilde{\mathcal{S}}_{k}$.
38:     for each local epoch $e\in[E,2E]$ do
39:         $\theta^{t,e+1}_{k}\leftarrow\theta^{t,e}_{k}-\eta_{\theta}\bigtriangledown\mathcal{L}_{3}$.
40:     end for
41:     $\theta^{t+1}_{k}=\theta^{t,2E}_{k}$.
42:end for
Algorithm 2 DualMinCE
1:
2:$\mathcal{D}_{p}$: the public dataset;
@@
4:$\mathcal{S}_{k}=\{(l_{k}^{i},p_{k}^{i})\}_{i=1}^{N},k=0\text{ or }\lceil K\rceil$: loss-logit pairs passed from either the server or clients.
5:$\mathcal{S}$.
6:$\tilde{\mathcal{S}}\leftarrow\{\}$triangleright initialize an empty set of selective knowledge.
7:for each $x^{i}$ in $\mathcal{D}_{p}$ do
8:     $l^{i}_{\text{local}}\leftarrow h(x^{i})$
9:     $k^{*}=\begin{cases}\arg\min\limits_{k}{(l^{i}_{k})},&\text{if }k=\lceil K\rceil\\ 0,&\text{if }k=0\end{cases}$
10:     $\tilde{\mathcal{S}}\leftarrow\tilde{\mathcal{S}}+(x^{i},p^{i}_{k^{*}})$ if $l^{i}_{k^{*}}<l^{i}_{\text{local}}$
11:end for
3.2 Bidirectional Token Alignment